### v2

@@ Abstract
In recent years, the advent of foundation models (FM) for digital pathology has relied heavily on scaling the pre-training datasets and the model size, yielding large and powerful models. While it resulted in improving the performance on diverse downstream tasks, it also introduced increased computational cost and inference time. In this work, we explore the distillation of a large foundation model into a smaller one, reducing the number of parameters by several orders of magnitude. Leveraging distillation techniques, our distilled model, H0-mini, achieves nearly comparable performance to large FMs at a significantly reduced inference cost. It is evaluated on several public benchmarks, achieving 3rd place on the HEST benchmark and 5th place on the EVA benchmark. Additionally, a robustness analysis conducted on the PLISM dataset demonstrates that our distilled model reaches excellent robustness to variations in staining and scanning conditions, significantly outperforming other state-of-the art models. This opens new perspectives to design lightweight and robust models for digital pathology, without compromising on performance.
@@ 1 Introduction
In recent years, representation learning models have revolutionized computational pathology (CPath) by providing efficient models that can serve for a variety of tasks, involving biomarker prediction [20, 28, 33], gene expression prediction [12, 13, 19, 30], whole slide image (WSI) and tissue classification [31, 39] or survival analysis [6]. Modern computational pathology frameworks typically rely on foundation models [1, 3, 14, 15, 23, 24, 29, 32, 34, 36, 41]. While differences exist between the various frameworks, they share the common idea to leverage representation models, able to map small patches of tissues (or tiles) to a lower dimensional space by computing features (or embeddings). Those features can then later be aggregated to produce predictions or representations at slide level.
@@
When designing CPath pipelines, two equally important questions arise: the model’s pure performance, and its robustness to variations in scanning conditions [16]. Recent studies have mostly focused on scaling the training dataset and the model size, resulting in large models whose size can reach more than one billion parameters, pretrained on several hundred thousands to a few million WSI [29, 36, 41]. This yielded spectacular improvements in performance on a variety of benchmarks and downstream tasks, yet, introduced an increased computational cost. Besides, the robustness of foundation models to variations in sample preparation (e.g., staining) or digitization (e.g., scanning) has not attracted the same attention as their performance despite being critical to the clinical deployment of CPath workflows. To further investigate this aspect, some datasets and benchmarks have been proposed, such as [9, 22, 25, 35].
@@
To jointly tackle the issue of the additional computational cost while promoting the robustness of the models, this study proposes to investigate the distillation of large foundation models. Starting from a recent foundation model, H-Optimus-0 [29], a Vision Transformer-giant (ViT-g) [10] with more than one billion parameters, we investigate its distillation into a smaller ViT-Base of 86 million parameters [10]. On several public benchmarks, the distilled model, H0-mini, demonstrates competitive performance with significantly larger state-of-the-art models. Additionally, leveraging the PLISM dataset [25], we show that the distilled model significantly outperforms all other foundation models reported in the literature for its robustness to variations in staining and scanning conditions. This uncovers new perspectives to design robust CPath models in view of their adoption in clinical practice. Our preprocessed version of PLISM dataset is publicly available for benchmark purposes at https://huggingface.co/datasets/owkin/plism-dataset-tiles.
2 Related work
2.1 Self-supervised learning (SSL) for digital pathology
Pretraining a feature extractor with SSL is now a cornerstone of modern CPath frameworks. First CPath pipelines used to leverage models pretrained on Imagenet [8], and suffered from an out-of-domain gap when transferred to digital pathology images. The advent of SSL methods showed the benefit of doing an in-domain pretraining for pathology feature extractors [5, 7, 27]. While it was originally leveraging contrastive methods [4, 38] tailored for convolutional networks, more recent models build on the Vision Transformer (ViT) architecture, with pretraining methods coming such as DINO, iBOT or DINOv2 [2, 26, 40]. Most of the recent foundation models for digital pathology leverage the DINOv2 framework [23, 29, 35, 36].
2.2 Distillation
@@
Distillation is a well known machine learning technique that consists in supervising a student model by a teacher model [18]. In the context of building image foundation models, it has been shown to be more efficient to distill a large model into a smaller one rather than training the small model from scratch [26]. In [11], a simple method is proposed to perform the distillation, and variations corresponding to the various SSL frameworks are investigated. In particular, following the DINO and iBOT frameworks, two methods are proposed, namely RoB-DINO and RoB-iBOT. In the former, following the DINO self-distillation objective, the class token of the teacher serves to supervise the student. In the latter, following the iBOT masked image modeling objective, an additional loss term is added where the (unmasked) patch tokens of the teacher also serve to supervise the student.
While distillation techniques are well established in computer vision in general, their application to foundation models for digital pathology remains relatively scarce. In GPFM [23], 3 foundation models (CONCH, Phikon and UNI) are simultaneously distilled during the pre-training phase of the model in addition to the classical DINOv2 loss, aiming to capture “experts” knowledge in the resulting model. In [41], a ViT-Small model, Virchow2G-Mini is introduced. It results from the distillation of Virchow2G (a ViT-Giant) on a 1-billion tile dataset. The authors show that distillation is beneficial compared to pre-training the same ViT-Small model from scratch.
@@ 3.1 Pre-training setup
3.1.1 Distillation setup
In this study, H-Optimus-0 is considered as the teacher model. We then follow the general methodology described in [26]. More precisely, for an image $x$, let $x_{1}$ and $x_{2}$ denote two augmented views of $x$. Additionally, for $i\in\llbracket 1,2\rrbracket$, we denote by $z_{i}^{(t)}$ (resp. $z_{i}^{(s)}$) the class tokens output by the teacher (resp. student) model for image $x_{i}$. For a patch $p$, we also denote by $z_{i,p}^{(t)}$ (resp. $z_{i,p}^{(s)}$) the patch tokens output by the teacher (resp. student) for the image $x_{i}$. We pass the teacher (resp. student) class tokens through the corresponding DINO head. This head is a MLP outputting a vector of scores, followed by softmax centering. Similarly, patch tokens are passed through either the teacher or student iBOT head. As a general notation, $h_{i}$ corresponds to the head projection of $z_{i}$. Taking inspiration from [11], distillation is performed by combining two objectives:
  • $\bullet$
    DINO objective. In this setting, only the class scores (derived from the projection of class tokens into the DINO heads) are used to perform the distillation. Let $H$ denote the cross-entropy loss. The corresponding loss function $L_{\text{dino}}$ is defined as
    $L_{\text{dino}}:=\bigg{(}H(h_{1}^{(t)},h_{2}^{(s)})+H(h_{2}^{(t)},h_{1}^{(s)})\bigg{)}/2.$
  • $\bullet$
    iBOT objective. The iBOT objective extends the distillation by incorporating patch tokens supervision. Only global crops are used to compute the objective. This loss is defined as the following
    $L_{\text{ibot}}:=\frac{1}{2P}\sum_{p=1}^{P}\sum_{j=1}^{2}H(h_{j,p}^{(t)},h_{j,p}^{(s)}),$
    where $P$ denotes the total number of patches. Unlike [11] we do not apply masking of the patches and the iBOT loss is applied for all patch scores (derived from the projection of patch tokens into the iBOT heads).
Finally, for all experiments, we keep a spare Exponential Moving Average (EMA) of the student as in [11] and we remove the stochastic depth. Separate heads for the DINO and iBOT distillation objectives are considered, and the Koleo regularizing loss is removed. To speed up the distillation, mixed precision is used as in [26]. We present in Table 4 the main parameters used for the distillation.
3.1.2 Pre-training datasets
To assess the benefit of distillation compared to pretraining feature extractors from scratch, and similar to the Phikon setup [14], we perform the distillation on a dataset of 43M tiles extracted from 6,093 TCGA slides covering 16 cancer sites. Compared to the typical sizes of pre-training datasets, it stands out by its relatively small size. We note that this is a difference with the pretraining setup proposed in [41], where the distillation is performed on a 1B-tile dataset.
@@ 3.2.1 Performance benchmarks
To rigorously evaluate model performance, we use two publicly available benchmarks [^1] The HEST benchmark is available at https://github.com/mahmoodlab/HEST, and the EVA benchmark is available at https://kaiko-ai.github.io/eva/main/.. The HEST-Benchmark [19] collects gene expression prediction tasks for 9 different indications. For each task, a subset of 50 highly variable genes is considered. To account for various embedding dimensions between feature extractors, we follow the recommended evaluation procedure from HEST-Benchmark, fitting a ridge regression on top of a PCA reduction with 256 components. We refer to the HEST paper and repository for more details on the evaluation procedure. Additionally, following the recommended methodology in [41], we use the concatenation of the class token and the mean over all patch tokens as embeddings of our distilled models. We note that because of the embedding dimension of ViT-Base models ($d$ = 768), concatenating the class token and the mean over all patch tokens results in embeddings of the same dimension as ViT-g models (e.g., H0 or Gigapath) ($d$ = 1536).
We also leverage the EVA benchmark [17], which consists in 4 patch-level classification tasks, 2 patch-level segmentation tasks, and 2 slide-level classification tasks. For each category of tasks, EVA’s evaluation protocol is fixed to allow for a fair evaluation across feature extractors. For instance, the main hyperparameters such as learning rate, batch size, or schedulers are shared for all feature extractors. We refer to the EVA paper and documentation for more details on the evaluation procedure.
Our results can be reproduced on HEST-Benchmark and EVA’s respective code repositories.
3.2.2 Robustness dataset
In order to evaluate model robustness, we use the public PLISM [25] dataset. PLISM consists of 46 human tissue types stained using 13 different H&E conditions, and captured using 7 WSI scanners. Of the 13 stainings, 3 were kept and 10 were discarded because they were deemed too different from routine H&E stainings. Each of those 21 experimental conditions consisted of one whole-slide image made of the 46 tissue punches. Those 21 WSIs were further aligned together using the Elastix software [21], taking the GMH-stained WSI captured with the Hamamatsu Nanozoomer S60 as reference. An in-house matter detection model was then used to detect tissue and tessellate those 21 WSIs into 16,278 tissue tiles each. Consequently, our robustness dataset is made of 16,278 tissue patches, each of them existing under 3 different H&E staining conditions and captured with 7 different scanners. We provide in Appendix 7.1 some visualizations from PLISM, following the registration step. Cosine similarity and top-k accuracy are used to evaluate the robustness of the feature extractors. These metrics are computed on slide pairs corresponding to different scanner and staining combinations.
  • $\bullet$
    Cosine similarity is assessed on features extracted from tissue tiles, averaged over all 16,278 matching tile pairs. We note that this metric was proposed in [35] to quantify the robustness of feature extractors to stain normalization.
  • $\bullet$
    Top-k accuracy computes the percentage of tiles from one slide whose matching tile on the other slide ranks among the k closest tiles (by cosine similarity) when compared to all other tiles from both slides. This metric is computed bidirectionally and averaged for each slide pair.
Finally, metrics are aggregated across unique combinations of distinct scanners and stainings, yielding a final robustness score for each feature extractor. We provide in Appendix 7.3 technical details on metrics computation. Specific robustness to scanner (resp. staining) is assessed on all intra-staining (resp. scanner) inter-scanner (resp. staining) pairs while overall robustness metrics is assessed using all pairs.
@@ 4 Results
4.1 HEST results
In Table 1, we report the results on the HEST benchmark. We notice that the distilled model performs remarkably well on this benchmark, highlighting the efficiency of knowledge distillation. The model significantly outperforms Phikon, its equivalent trained from scratch (same architecture and same pre-training data). Additionally, both versions outperform in average much bigger foundation models such as Virchow2 or Prov-Gigapath.
Table 1: HEST results. Pearson correlations are reported for all tasks. For all models, [CLS] tokens are concatenated to the mean over all patch tokens to form the input embeddings.1 For UNI 1.5, results are directly taken from HEST repository’s README.
Model Size IDC PRAD PAAD SKCM COAD READ ccRCC LUAD IDC Average
H-Optimus-0 1,100M 0.6106 0.3621 0.5106 0.6614 0.3089 0.2401 0.2669 0.5754 0.2664 0.4224
UNI 1.51 - 0.5989 0.3645 0.4902 0.6401 0.2925 0.2240 0.2522 0.5586 0.2695 0.4090
H0-mini 86M 0.5909 0.3633 0.5068 0.6125 0.2700 0.2047 0.2643 0.5633 0.2640 0.4044
Virchow2 632M 0.5971 0.3528 0.4778 0.6404 0.2580 0.2073 0.2604 0.5685 0.2568 0.4019
Hibou Large 307M 0.5945 0.3231 0.4758 0.6059 0.3128 0.1823 0.2777 0.5720 0.2490 0.3992
Kaiko ViT-B/8 86M 0.5710 0.3827 0.4727 0.5904 0.3105 0.1726 0.2664 0.5883 0.2362 0.3912
UNI 307M 0.5851 0.3274 0.4882 0.6235 0.2583 0.1757 0.2463 0.5558 0.2576 0.3907
Prov-Gigapath 1,100M 0.5707 0.3841 0.4920 0.5823 0.3076 0.186 0.2277 0.5579 0.2499 0.3952
Phikon-V2 307M 0.5677 0.3793 0.4771 0.5845 0.2561 0.1865 0.2607 0.5502 0.2476 0.3897
Phikon 86M 0.5481 0.3452 0.4639 0.5555 0.2668 0.1667 0.2496 0.5679 0.2387 0.3780
4.2 EVA results
In Table 2, we report the results on the EVA benchmark. As suggested by the authors of the benchmark, we report the average results with and without the BACH task, as the spatial resolution of the images differ from the other tasks, and therefore tend to favor mixed-magnification models such as Virchow2 or the Kaiko models. Even though the results tend to saturate, with 6 models with an average performance (without BACH) between 0.78 and 0.79, similar conclusions from the HEST benchmark can be drawn. The distilled models are competitive with the other state-of-the-art models while being much smaller. A noticeable exception is the Kaiko-B/8 model which, with a similar architecture (but a smaller patch size which results in more computations), reaches equivalent performance. This shows that the distilled model can perform very well for a variety of tasks, including patch classification, WSI classification (ranking first) or segmentation.
Table 2: EVA results. Balanced binary (resp. multiclass) accuracy is reported for binary (resp. multiclass) classification tasks. MonaiDiceScore is reported for segmentation tasks. For all models, input embeddings are only the [CLS] tokens. Following EVA benchmark, average performance is taken over all tasks except BACH.
Patch-level classification Slide-level classification Segmentation Average

### v3

@@ Abstract
In recent years, the advent of foundation models (FM) for digital pathology has relied heavily on scaling the pre-training datasets and the model size, yielding large and powerful models. While it resulted in improving the performance on diverse downstream tasks, it also introduced increased computational cost and inference time. In this work, we explore the distillation of a large foundation model into a smaller one, reducing the number of parameters by several orders of magnitude. Leveraging distillation techniques, our distilled model, H0-mini, achieves comparable performance to large FMs at a significantly reduced inference cost on HEST and EVA public benchmarks. Additionally, we conduct robustness analyses on the PLISM dataset and a multi-scanner, multi-staining private breast cancer cohort. We demonstrate that our distilled model reaches excellent robustness to variations in staining and scanning conditions, significantly outperforming other state-of-the art models. This opens new perspectives to design lightweight and robust models for digital pathology, without compromising on performance. We publicly release H0-mini [^1] https://huggingface.co/bioptimus/H0-mini along with plismbench [^2] Available at https://github.com/owkin/plism-benchmark., the first robustness benchmark of pathology foundation models based on the PLISM dataset.
@@ 1 Introduction
In recent years, representation learning has revolutionized computational pathology (CPath) by providing efficient models that can serve for a variety of tasks, including biomarker prediction, gene expression prediction, whole slide image (WSI) and tissue classification or survival analysis [22]. Modern computational pathology frameworks typically rely on foundation models [3, 6, 7, 17, 18, 21, 24, 27], and share the common idea to leverage representation models (or feature extractors), able to map small patches of tissues (or tiles) to a lower dimensional space by computing features (or embeddings). Due to label scarcity at the tile-level, recent pathology feature extractors have been pretrained using self-supervised learning (SSL) methods such as DINO, iBOT or DINOv2 [2, 20, 26] and have become a cornerstone of modern CPath frameworks. Recent studies have mostly focused on scaling the pretraining dataset and the model size, resulting in large models called foundation models (FMs) whose size can reach more than one billion parameters, pretrained in up to a few million WSIs using the DINOv2 framework [21, 24, 27]. This yielded spectacular improvements in performance on a variety of downstream tasks, yet introduced an increased computational cost. Besides, the robustness of FMs to variations in sample preparation or digitization has not attracted the same attention as their performance, which is critical to the clinical deployment of CPath workflows [8]. Some datasets and benchmarks have been proposed to investigate this aspect [15, 19, 23].
@@
To jointly tackle the issue of the additional computational cost while promoting the robustness of the models, this study proposes to investigate the distillation of large FMs. Starting from a recent foundation model, H-Optimus-0 [21], a Vision Transformer (ViT)-giant [4] with more than one billion parameters, we investigate its distillation into a smaller ViT-Base of 86 million parameters. On several public benchmarks, the distilled model, H0-mini, demonstrates competitive performance with significantly larger state-of-the-art models. Additionally, leveraging the PLISM dataset [19] and a private multi-scanner, multi-staining dataset for breast cancer biomarkers prediction, we show that the distilled model significantly outperforms all other foundation models reported in the literature in its robustness to variations in staining and scanning conditions. This uncovers new perspectives to design robust CPath models in view of their adoption in clinical practice.
1.0.1 Related work.
@@
Distillation is a well known machine learning technique that consists in supervising a student model by a teacher model [10]. In the context of building image foundation models, it has been shown to be more efficient to distill a large model into a smaller one rather than training the small model from scratch [20]. In [5], a simple method is proposed to perform the distillation, and variations corresponding to the various SSL frameworks are investigated such as RoB-DINO and RoB-iBOT. While distillation techniques are well established in computer vision in general, their application to FMs for digital pathology remains relatively scarce. In GPFM [17], 3 foundation models (CONCH [16], Phikon and UNI) are simultaneously distilled into a student model in addition to DINOv2 losses. In [27], a ViT-Small model Virchow2G-Mini is introduced, resulting from the distillation of Virchow2G (a ViT-Giant) on one billion tiles.
@@ 2.1 Pre-training setup
2.1.1 Distillation setup.
In this study, H-Optimus-0 (H0, [21]) is considered as the teacher model. We then follow the general methodology described in [20]. More precisely, for an image $x$, let $x_{1}$ and $x_{2}$ denote two augmented views of $x$. Additionally, for $i\in\llbracket 1,2\rrbracket$, we denote by $z_{i}^{(t)}$ (resp. $z_{i}^{(s)}$) the class tokens output by the teacher (resp. student) model for image $x_{i}$. For $i\in\llbracket 1,2\rrbracket$ and for a patch $p$, we also denote by $z_{i,p}^{(t)}$ (resp. $z_{i,p}^{(s)}$) the patch tokens output by the teacher (resp. student) for the image $x_{i}$. We pass the teacher and student class tokens (resp. patch tokens) through the corresponding DINO (resp. iBOT) head. $h_{i}$ denotes the vector of prototype scores output from the head projection of $z_{i}$. Taking inspiration from [5], distillation is performed by combining two objectives.
DINO objective.
In this setting, only the class scores are used to perform the distillation. Let $H$ denote the cross-entropy loss. The corresponding loss function $L_{\text{dino}}$ is defined as
$L_{\text{dino}}:=\bigg{(}H(h_{1}^{(t)},h_{2}^{(s)})+H(h_{2}^{(t)},h_{1}^{(s)})\bigg{)}/2.$ (1)
IBOT objective.
The iBOT objective extends the distillation by incorporating patch scores supervision. This loss is defined as the following
$L_{\text{ibot}}:=\frac{1}{2P}\sum_{p=1}^{P}\sum_{j=1}^{2}H(h_{j,p}^{(t)},h_{j,p}^{(s)}),$ (2)
where $P$ denotes the total number of patches. Unlike [5] we do not apply masking of the patches and the iBOT loss is applied for all patch scores.
Finally, for all experiments, we keep a spare exponential moving average (EMA) of the student as in [5], we remove the stochastic depth and Koleo regularization. To speed up the distillation, mixed precision is used as in [20]. H0-mini was trained for 105,000 iterations and a batch size of 2,048 using 128 Nvidia V100 32Go for a total of 4,350 gpu hours. For all downstream experiments, we used the spare EMA of the student as the feature extractor. The main hyperparameters used in the distillation are given in appendix 0.A.1.
2.1.2 Pre-training datasets.
We perform the distillation on a dataset of 43M tiles (224x224 at 20x magnification) extracted from 6,093 TCGA slides covering 16 cancer sites. This allows for a fair comparison with Phikon, a ViT-B feature extractor pretrained from scratch on the same dataset with iBOT. Compared to the typical sizes of pre-training datasets, it stands out by its relatively small size. In particular, we note the difference with the pretraining setup proposed in [27], where the distillation is performed on a 1B-tile dataset.
@@ 2.2 Evaluation setup
To rigorously evaluate model performance, we use two publicly available benchmarks [^3] The HEST benchmark is available at https://github.com/mahmoodlab/HEST, and the EVA benchmark is available at https://kaiko-ai.github.io/eva/main/.. The EVA benchmark [9] consists in 4 patch-level classification tasks, 2 patch-level segmentation tasks, and 2 slide-level classification tasks. For each category of tasks, EVA’s evaluation protocol is fixed to allow for a fair evaluation across feature extractors (e.g., learning rate, batch size or schedulers). We also leverage the HEST-Benchmark [12] which collects gene expression prediction tasks for 9 different indications. For each task, a subset of 50 highly variable genes is considered. To account for various embedding dimensions between feature extractors, we follow the recommended evaluation procedure from HEST-Benchmark, fitting a ridge regression on top of a PCA reduction with 256 components. Additionally, and following the recommended methodology in [27], we use the concatenation of the class token and the mean over all patch tokens as embeddings of our distilled model.
2.2.2 PLISM robustness dataset.
To evaluate model robustness, we first use the public PLISM [19] dataset which consists of 46 human tissue types stained using 13 different H&E conditions, and captured using 7 WSI scanners (Figure 1). This results in 91 WSIs later registered using the Elastix software [14] and tessellated into 16,278 tissue tiles each. We make this processed dataset publicly available [^4] The processed PLISM dataset is made available at https://huggingface.co/datasets/owkin/plism-dataset.
Cosine similarity and top-$k$ accuracy are used to evaluate the robustness of the feature extractors across scanner and staining combinations. Cosine similarity is assessed on features extracted from tissue tiles averaged over all matching tile pairs. We note that this metric was proposed in [23] to quantify the robustness of feature extractors to stain normalization. Top-k accuracy computes the percentage of tiles from one slide whose matching tile on the other slide ranks among the $k$ closest tiles (by cosine similarity) when compared to all other tiles from both slides. Finally, metrics are aggregated across slide pairs. Specific robustness to scanner (resp. staining) is assessed on all fixed-staining (resp. scanner) cross-scanner (resp. staining) pairs. We also report overall robustness on cross-stainings, cross-scanners slide pairs. Additional details on the implementation of those metrics are given in appendix 0.A.2.
2.2.3 BreastBiomarker (BreastBm) performance & robustness benchmark.
We also evaluate the downstream robustness of H0-Mini to variations in scanner and staining in the context of breast cancer biomarker prediction. Slide-level tasks aim at predicting the positivity of estrogen receptor status (ER), progesterone receptor status (PR), human epidermal growth factor receptor 2 status (HER2) and germline BRCA1/BRCA2 mutations (gBRCA). We leverage 3 private cohorts for robustness evaluation (see Figure 1). The “GR” (resp. “CLB”) cohorts consists of 671 (resp. 353) tumor blocks providing 2 consecutive slides: one stained with Hematoxylin Eosin Saffron (HES) (resp. archival Hematoxylin Phloxine Saffron, HPS) and one with H&E. All GR (resp. CLB) slides are scanned using an Olympus VS200 (resp. Aperio GT450 and Aperio AT2). GR H&E slides are also scanned with Aperio ScanScope and Hamamatsu S60. This yields 6 possible subcohort pairs for GR and for CLB. The CuriMeta cohort consists of 208 H&E slides scanned with Pramana and Huron scanners (1 pair).
For all endpoints and to best isolate the impact of feature extractors, all downstream models use a mean pooling architecture: a logistic regression is trained on the average of tile-level feature representations. Performance is assessed using AUC while robustness is evaluated via the Concordance Correlation Coefficient (CCC) between WSI-level predictions from different stainings and scanners for the same tumor blocks. For ER and PR (resp. HER2) endpoints, training was performed on 1,109 WSIs from TCGA-BRCA and evaluation on GR, CLB and CuriMeta cohorts (resp. GR cohort). Due to labels unavailability in TCGA, predictive models for the gBRCA endpoint were trained on GR and validated on CLB and CuriMeta.
Figure 1: Visualizations of BreastBm (left) and PLISM (right) datasets. For PLISM, we only display 7 of the 13 different stainings on the y-axis.
@@ 3 Results
3.1 Performance evaluation
In Table 1, we report the results on the EVA benchmark. As suggested by its authors, we report the average results without the BACH task, as the spatial resolution of the images differs from the other tasks and therefore tend to favor mixed-magnification models such as Virchow2 or the Kaiko models [1]. Even though the results tend to saturate (6 models show an average performance between 0.78 and 0.79), H0-mini is competitive with the other state-of-the-art models while being much smaller. A noticeable exception is the Kaiko-B/8 model reaching equivalent performance (but with a smaller patch size which results in more computations). This shows that the distilled model can perform well on a variety of tasks, including patch classification, WSI classification or segmentation.
Table 1: EVA results. For all models, input embeddings are only the [CLS] tokens. Balanced accuracy (resp. “MonaiDiceScore”) is reported for classification (resp. segmentation) tasks. () results may be over-estimated due to an overlap between downstream and pre-training datasets.
Patch-level classification Slide-level classification Segmentation Mean